OpenStudy (anonymous):

The squares of two consecutive negative integers total 85.Find the lesser of the numbers.

OpenStudy (javk):

let the lower of the two integers= x therefore the larger integer =(x+1) because the two numbers are consecutive add the two squares of these two together \[x ^{2}+ (x+1)^{2}=85\] expand the brackets \[x ^{2} + x ^{2}+2x+1=85\] take everything to the right hand side and neaten \[2x ^{2}+2x-84=0\] you can simplify the equation by dividing everything by 2 \[x ^{2}+x-42\] now solve the eqn.

OpenStudy (yttrium):

Let: Smaller integer = -x Larger integer = -x+1 (-x+1)^ + (-x)^2 = 85 x^2 - 2x + 1 + x^2 = 85 2x^2 - 2x -84 = 0 x^2 - x - 42 = 0 x = -7, 6 Therefore, the lesser of the numbers is: -x = -(-7) = 7 (not valid) -x = -(6) = -6 valid (Quite Easily Done)

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

The difference of two positive numbers is 3 and the sum of their squares is 225.What are the numbers

OpenStudy (anonymous):

sorry I dont get that

OpenStudy (javk):

If you got what i did before, do the same thing a gain Step 1: give your variables a name lets use 'y' this time let the lower of the two integers = y therefore the larger integer = (y+3) Step 2: Figure out what the eqn would look like \[y ^{2}+ (y+3)^{2}=225\] Step 3: Bring it into a format that will be easier for you to solve -expand the brackets \[y ^{2} + y ^{2}+6y+9=225\] -take everything to the right hand side and neaten \[2y ^{2}+6y-216=0\] you can simplify the equation by dividing everything by 2 \[y ^{2}+3y-108\] Step 4:Now solve the eqn. -factorize the eqn \[(y+12)(y-9)=0\] either \[y+12=0\], \[y=-12\] -or \[y-9=0\], \[y=9\] since it says in the ques they have to be positive values, reject \[y=-12\] therefore the lower of the two no.s is 9 consequently making the other number 12